# ToDo (#1): #2

### v3

@@
Abstract: This Preliminary Design Report (PDR) describes the IsoDAR electron-antineutrino source in two volumes which are mostly site-independent and describe the cyclotron driver, providing a 10 mA proton beam, and the medium energy beam transport line and target, respectively. The IsoDAR driver and target will produce a mole of electron-antineutrinos over the course of five years. Paired with a kton-scale liquid scintillator detector, it will enable an impressive particle physics program including searches for new symmetries, new interactions and new particles. Here in Volume I, we describe the driver, which includes the ion source, low energy beam transport, and cyclotron. The latter features radiofrequency quadrupole (RFQ) direct axial injection and represents the first accelerator purpose-built to make use of so-called vortex motion.
Section Editors (in alphabetical order): M. Abs1, J.R Alonso2, J.M. Conrad2, S.J.B. Engebretson2, E. Forton1, A. Herrod1, D. Joassin1, J. Moon2, S. de Neuter1, E. Van der Kraaij1, G. Wéry1, D. Winklehner2, E. Winkler2
for the IsoDAR collaboration
1Ion Beam Applications, 3 chemin du cyclotron Louvain-la-Neuve, BELGIUM
2Massachusetts Institute of Technology, 77 Massachusetts Ave. Cambridge, MA, USA
Corresponding Authors: Joshua Spitz (spitzj@umich.edu) and Daniel Winklehner (winklehn@mit.edu)
V2.0
Chapter 1 Introduction
Figure 1.1: Schematic of the IsoDAR experiment installed at the Yemilab facility in South Korea. From right to left: The compact cyclotron with RFQ injector, the medium energy beam transport, the target, and the LSC neutrino detector. From Ref. [calvo_analyzing_2023].
IsoDAR will be the first high-intensity, proton-driven source to be located alongside an underground multi-kiloton-scale scintillator detector. The first version of IsoDAR will be constructed at the Yemilab facility [alonso_isodaryemilab_2022] within the Handuk mine, located in South Korea where the “Liquid Scintillator Counter” (LSC) [^1] The LSC was renamed $\nu$Eye in March 2025, but for consistency throughout this document, we will refer to the detector as the LSC. is planned to be installed [seo_physics_2023]. The novel source, consisting of a cyclotron, beam transport line, and a target/sleeve is shown in Fig. 1.1. In five calendar years (four live years), the source will produce 1.67 million $\bar{\nu}_{e}+p\rightarrow e^{+}+n$ (inverse beta decay or ”IBD”) events and 7000 $\bar{\nu}_{e}+e^{-}\rightarrow\bar{\nu}_{e}+e^{-}$ (elastic scattering or ES) events in the LSC that can be used for searches for new physics. More than a mole of neutrons and mono-energetic photons are also produced in the target, and these may act as progenitors for new particles that enter the LSC and leave signatures from interactions and decays. IsoDAR can address a broad range of new physics manifested as new symmetries, new interactions and new particles that is inaccessible to traditional experiments. For further discussion see Refs. [alonso_neutrino_2022, waites_axionlike_2023].
@@
This document is Volume I in a two-volume Preliminary Design Report (PDR) for IsoDAR that provides the engineering solutions necessary to establish credible solutions to the key issues and cost-drivers identified during the past decade of conceptual design and R&D. The contents of the PDR will cover the Technical Facility: the cyclotron, the beam transport and the target/sleeve. This document builds on the structure and text of the Conceptual Design Report (CDR) of the IsoDAR Technical Facility [abs_isodarkamland_2015]. This step of producing a PDR substantially reduces risk related to IsoDAR approval: the remaining detailed engineering that will follow will build on these results and be straightforward to estimate in scope and cost.
@@
Volume I describes the IsoDAR proton driver. This consists of a “front end” that delivers ions to a cyclotron for acceleration and extraction and the cyclotron itself. This document presupposes a basic understanding of cyclotrons; for an introduction to cyclotrons and a cost/benefit analysis of this choice of machine for IsoDAR, see the discussion in the CDR.
@@
The 60 MeV/amu IsoDAR Proton Driver is exceptional in that it is designed and engineered to produce an order of magnitude higher current proton beams at extraction than commercial machines of similar energy. Achieving this high power necessitated three key breakthroughs. First, the cyclotron accelerates $\mathrm{H}_{2}^{+}$, mitigating “space charge effects” (electromagnetic interactions between beam particles). Second, the beam is axially injected through an RFQ, acting as a high-efficiency buncher. Third, the design harnesses vortex motion, a complex effect previously observed at PSI Injector II [stetson:vortex, baumgarten:vortex1], inducing stable spiraling motion in the high-current accelerated beam, significantly reducing beam growth. For more concerning the accelerator physics that shapes the IsoDAR design, see Ref. [winklehner_order--magnitude_2022, winklehner_high_2018].
@@
These three physics requirements for reaching high current drive many of the engineering choices presented in this PDR. The other design driver is the need to construct the cyclotron underground. The cavern for the IsoDAR Proton Driver has already been constructed at Yemilab, and so the dimensions and constraints are understood, leading to fully realistic engineering plans.
@@
  • Chapter 1: (This chapter) proceeds with an overview of the physics searches that will be possible with IsoDAR, followed by an overview of the requirements for the proton driver.
  • Chapter 2: Describes engineering related to the IsoDAR front end.
  • Chapter 3: Describes engineering related to the general IsoDAR cyclotron design.
  • Chapter 4: Describes specifics of the design related to installation underground.
1.1 The IsoDAR Physics Case
Figure 1.2: Energy dependence of fluxes. Left: $\bar{\nu}_{e}$, unit normalized; Middle: photons per proton on target (POT) from excited nuclei [alonso_neutrino_2022]. Right: kinetic energy of neutrons at production in target and sleeve, normalized per proton.
IsoDAR offers a rich BSM physics program if paired with a kton-scale liquid scintillator detector like the proposed LSC at Yemilab [seo_physics_2023]. The exceptional reach is owed to the high flux rates of $\bar{\nu}_{e}$, photons, and neutrons from the $\approx 2\cdot 10^{24}$ protons on target over 5 years (cf. Fig. 1.2), and placing the source very close to the detector. Here we briefly discuss the highlights of the IsoDAR BSM physics program enabled by a 2.26 kton detector like the LSC in three categories: “New Particles” - experimentally and theoretically motivated; “New Symmetries” - CP-violation and the QCD Axion; And “New Interactions” - Precision Electroweak Physics. See Refs. [alonso_neutrino_2022, Hostert:2022ntu, waites_axionlike_2023] for further details.
1.1.1 Fluxes in IsoDAR
We consider three fluxes of particles produced in the IsoDAR target for BSM physics searches: the $\bar{\nu}_{e}$ flux, the neutron flux, and the $\gamma$ flux from the excitation of nuclei as neutrons are moderated. These fluxes are shown in Fig. 1.2.
The $\bar{\nu}_{e}$ flux stems from a single beta-decay of 8Li, which is produced by neutron capture on 7Li and is well-understood. It peaks at 6 MeV, which puts the majority of the events above backgrounds from the environment. These are clear advantages over reactor-neutrino fluxes. Interactions in the detector at these neutrino energies are exclusively inverse-beta-decay (IBD) and elastic scattering (ES). Cross sections for these reactions are known with uncertainties at $<1$% at these energies.
Neutrons in the IsoDAR target are primarily produced by spallation interactions of the beam protons with the beryllium target hemispheres. Additional secondary neutrons stem from interactions of the primary neutrons with beryllium in the target and sleeve. We model the processes in the target with Geant4 using an accurate representation of the latest target and sleeve design (see Refs. [bungau_optimizing_2019]). We show the resulting neutron energy spectrum in Fig 1.2 (right). Most neutrons capture on the 7Li in the sleeve. To ensure $<5$ neutrons per year enter the LSC veto region, the target is further shielded using boron-loaded concrete and steel [bungau_shielding_2020].
The third flux we consider is that of photons as seen in Fig.1.2 (middle). Here, one can observe a line structure that stems from the transitions of excited nuclei and and underlying smooth continuum due to bremsstrahlung. Again, the spectrum was generated using Geant4 using the full IsoDAR Target geometry [waites_axionlike_2023].
1.1.2 New Particles: Experimentally and Theoretically Motivated
Figure 1.3: An example of a potential 3+2 signal in IsoDAR, orange–true signature, black–IsoDAR reconstruction with uncertainty shown.

### v4

@@ Abstract
This Preliminary Design Report (PDR) describes the IsoDAR electron-antineutrino source in two volumes which are mostly site-independent and describe the cyclotron driver providing a 60 MeV, 10 mA proton beam (this Volume); and the medium energy beam transport line (MEBT) and target (Volume II). The IsoDAR driver and target will produce about $1.15\cdot 10^{23}$ electron-antineutrinos over five years. Paired with a kton-scale liquid scintillator detector, it will enable a broad particle physics program including searches for new symmetries, new interactions and new particles. Here in Volume I, we describe the driver, which includes the ion source, low energy beam transport, and cyclotron. The latter features Radio-Frequency Quadrupole (RFQ) direct axial injection and represents the first accelerator purpose-built to make use of so-called vortex motion.
keywords:
Neutrinos, High-Intensity, Cyclotron, BSM Physics
1 Introduction
Figure 1: Main: The IsoDAR accelerator and target system adjacent to the LSC detector at Yemilab. Top: A to-scale rendering of the cyclotron overlaid on a photo of the cavern. The project comprises four major systems, three of which are shown: 1) the cyclotron; 2) the 60 MeV beam transport; and 3) the target and sleeve. The fourth is a data acquisition and monitoring system for all components.
[…]
@@ 1 Introduction
This document is Volume I in a two-volume Preliminary Design Report (PDR) for IsoDAR that provides the engineering details necessary to establish credible solutions to the key issues and cost-drivers identified during the past decade of conceptual design and R&D. The contents of the PDR will cover the Technical Facility: the cyclotron, the beam transport and the target/sleeve. This document builds on the structure and text of the Conceptual Design Report (CDR) of the IsoDAR Technical Facility [3]. This step of producing a PDR substantially reduces the risk related to IsoDAR approval. The remaining detailed engineering that will follow will build on these results and will be straightforward to estimate in terms of scope and cost.
@@
This document presupposes a basic understanding of cyclotrons. For an introduction to cyclotrons and a cost/benefit analysis of this choice of machine for IsoDAR, see the discussion in the CDR.
@@
The 60 MeV/amu IsoDAR Proton Driver is exceptional in that it is designed and engineered to produce an order of magnitude higher current proton beams at extraction than commercial machines of similar energy. Three novel concepts set this accelerator apart from commercial ones: First, the cyclotron accelerates $\mathrm{H}_{2}^{+}$, mitigating “space charge effects” (electromagnetic interactions between beam particles). Second, the beam is axially injected through an RFQ, acting as a high-efficiency buncher. Third, the design harnesses “vortex motion”, a complex effect previously observed at the Paul Scherrer Institute at the PSI Injector II cyclotron [4, 5]. Vortex motion induces spiraling within the individual beam bunches, providing compression and stability, which leads to substantially lower losses during extraction. For more information concerning the accelerator physics that shapes the IsoDAR design, see Refs. [6, 7].
@@
These three concepts drive many of the engineering choices presented in this PDR. The other design driver is the need to assemble the cyclotron underground. The cavern for the IsoDAR source has already been constructed at Yemilab, and so the dimensions and constraints are understood, leading to realistic engineering plans.
@@
  • Chapter 2: Proceeds with an overview of the physics searches that will be possible with IsoDAR.
  • Chapter 3: Presents a brief overview of the requirements for the proton driver.
  • Chapter 4: Describes engineering related to the IsoDAR front end.
  • Chapter 5: Describes engineering related to the general IsoDAR cyclotron design.
  • Chapter 6: Describes specifics of the design related to installation underground.
2 IsoDAR Physics That Informs Design Goals
The physics case for IsoDAR has driven the design decisions that are reported in this Preliminary Design Report. In this section, we provide a relatively high-level overview of the physics goals and connect them to the design needs, providing references for the reader to pursue more details.
Today, particle physicists build massive underground detectors to capture low-energy interactions with unparalleled precision, while accelerator physicists develop powerful machines for new physics above ground; but these instruments have never been joined in close proximity. IsoDAR aims to change that paradigm. It will be the first high-intensity accelerator located underground, next to a multi-kiloton-scale detector, opening entirely new windows for particle physics.
The designs presented in this document are largely site-independent. Still, we will use details of the planned initial run at Yemilab in Korea to provide a concrete description of the physics impact. Figure 1 shows a 3D rendering of IsoDAR’s planned installation at Yemilab. The volumes of this PDR focus on the IsoDAR source, which consists of the elements to the right of the LSC (cylinder at left). More information on the LSC is available in Refs. [2, 8, 9]. Important quantities within this layout that impact the physics are listed in Table 1. Although this first volume concentrates only on the accelerator, the physics case relies on all systems, and this chapter will also be relevant to Volume 2 of the PDR.
Table 1: Specifications for the IsoDAR@Yemilab design that directly impact the physics opportunities. See text for explanations of the live-time, fiducial volume cut and energy minimum cut.
Accelerator Run Parameters
Runtime 5 years
Duty factor 80%
Livetime 4 years
Protons on IsoDAR target/year $1.97\cdot 10^{24}$
Expected total antineutrino flux in 4 yr livetime
$\bar{\nu}_{e}$ $1.15\cdot 10^{23}$
@@ 2 IsoDAR Physics That Informs Design Goals
LSC design parameters
LSC fiducial mass 2.26 kton
Energy resolution at 3 MeV 3.7%
Baseline distance from source 17 m center-to-center (9.5-25.6 m)
Depth of LSC and source 2700 m.w.e.
SM rates with livetime, energy and fiducial volume assumptions
IBD ($\bar{\nu}_{e}+p\rightarrow e^{+}+n$) 1.67 million
ES ($\bar{\nu}_{e}+e^{-}\rightarrow\bar{\nu}_{e}+e^{-}$) 7000
2.1 Introduction to Design for BSM Searches in IsoDAR
One typically searches for Beyond Standard Model (BSM) physics through its effect on Standard Model (SM) particles in three ways. First, through the production of a new particle via coupling to an SM particle that subsequently interacts or decays. Second, through the observation of new symmetries that introduce unexpected behavior of SM particles. Third, modifications to SM interactions due to new forces that couple to SM particles. From this, one sees that a crucial key to the discovery of new physics lies in the SM flux available to an experiment. The higher the quality and quantity of the SM flux, the deeper the reach in the search for new physics.
The IsoDAR system focusses 60 MeV protons onto a beryllium target that is surrounded by a 99.99% pure 7Li sleeve that, in turn, is embedded in substantial shielding. Given this design, IsoDAR produces three unique SM fluxes for BSM searches: electron antineutrinos, monoenergetic photons and neutrons. We discuss the qualities and quantities of these fluxes below, and they are shown in Fig. 2. These three SM fluxes fall into two categories of BSM physics that can be explored. The first category consists of SM particles that can exit the IsoDAR source and cross the shielding, ultimately producing signatures in the detector — electron antineutrinos. The second category is the set of copiously produced SM particles — the photons and neutrons that interact electromagnetically and strongly, respectively. A new BSM particle can couple to these contained fluxes and, if it has a very low interaction cross-section, will traverse the shielding and enter the detector. To be observed, the particle must subsequently mix with, interact with, or decay to SM particles that are then observable in the LSC. The IsoDAR design has three crucial advantages for BSM searches using these particles:
  1. 1.
    The rate of SM particle production is very high.
  2. 2.
    The solid angle to the LSC detector, 17 m from the source, is very large.
  3. 3.
    The cosmogenic and environmental backgrounds are very low.
We consider the importance of each of these points below.
Figure 2: Energy dependence of fluxes. Left: $\bar{\nu}_{e}$, normalized to unity; Middle: photons per Proton On Target (POT) from excited nuclei. Right: kinetic energy of neutrons at production in the target and sleeve, normalized per proton.
2.2 Designing IsoDAR’s Electron Antineutrino Flux
IsoDAR was originally designed for BSM searches using the electron-antineutrino flux. Although the physics program has since expanded to include state-of-the-art searches using both photon and neutron fluxes, the collaboration has prioritized the quality of the electron-antineutrino flux as a primary design goal. The quality issues that have been maximized by the design are: understanding the spatial dependence of the flux, understanding the energy dependence of the flux, and purity of antineutrino content (i.e., minimizing backgrounds). The quantity of the electron-antineutrino flux is also a design priority, where rates ahve been balanced against cost and practicality of installing and running at an underground site. The outcome of the design is a world-leading rate of electron-antineutrino interactions, which we quote for a live time of 4 years. IsoDAR will collect 1.67 million $\bar{\nu}_{e}+p\rightarrow e^{+}+n$ (inverse beta decay or “IBD”) events and 7000 $\bar{\nu}_{e}+e^{-}\rightarrow\bar{\nu}_{e}+e^{-}$ (elastic scattering or “ES”) events in five calendar years—unprecedented sample sizes to probe new interactions [10]. These are the only two interactions open within the SM for $\bar{\nu}_{e}$ scattering at $\sim 6$ MeV. The IBD signature requires a coincidence between the initial positron light-flash and the subsequent neutron capture, which greatly reduces backgrounds compared to the single-flash ES signature. To further minimize background, both the positron (for IBD) and electron (for ES), signals are required to be reconstructed to >3>3 MeV and to be within the 2.26 kt (1.15 kt) fiducial volume for IBD (ES) events. The resulting backgrounds to the IBD signal are found to be negligible, while the assumed backgrounds for physics from the ES signal, with all cuts, are shown in Fig. 3.
Figure 3: The ES signal and background event rates expected in 5 years of running IsoDAR@Yemilab. No energy smearing has been applied to the distributions. From Ref. [10].
Some of the neutrino physics goals require precise knowledge of the spatial distribution of the beam. To this end, an isotope decay-at-rest (hence the name IsoDAR) flux was selected. This flux has the advantage of being produced in $4\pi$ steradians, and the relative flux within a given solid angle is straightforward to determine.
An isotope that $\beta$-decays with high endpoint (or Q-value) is desirable since the interaction rate in the detector increases with energy. After studying many isotopes, the collaboration has chosen 8Li, which has the well-predicted $\beta$ decay spectrum [11], shown in Fig. 2, that peaks at approximately 6 MeV and has an endpoint of 16 MeV, minimizing systematic uncertainties on the energy dependence of the flux. By choosing a flux with high endpoint, we can apply a visible energy cut above 3 MeV, removing most environmental background events while retaining 10% of the total flux.
The tradeoff is that a high Q-value corresponds to a short beta decay lifetime—839 ms for 8Li—requiring continuous replenishment via an accelerator-based system that produces neutrons that capture on 7Li. This leads to several design choices for the target, the surrounding “sleeve” and the accelerator that we discuss below.
The design requires a target that generates high quantities of neutrons at energies that must then be moderated efficiently to allow capture. After evaluating alternatives and considering radiation handling issues, we selected 9Be as the target, since its final neutron is bound by only 1.7 eV. A tungsten target would release more neutrons per proton on target, but would also generate unwanted neutrinos in the flux. An important design choice was to cool the target using a 17-liter closed-loop system of D2O, as opposed to H2O, which has a high neutron capture cross section.
The neutrons exit the target and enter a surrounding region with 7Li, which we refer to as the “sleeve.” This sleeve serves three purposes: to multiply neutrons, to moderate fast neutrons, and to provide the 7Li target for capture. Natural lithium is a mixture of 6Li and 7Li. Due to the very high 6Li neutron capture cross section, we require isotopic purification to 99.99% 7Li. Accessing 480 kg of 7Li at this purity has required coordination with industry. A cost-effective method to further enhance neutron production is to add beryllium to the sleeve. As the fast neutrons moderate, they inelastically scatter off the beryllium, releasing additional neutrons. We found the 8Li production rate is maximized with a 25% 7Li / 75% 9Be mixture. The goal is to maximize neutron absorption in 7Li while minimizing losses in all other elements of the target and sleeve and minimize background neutrino flux rates, hence the design focuses on low-Z materials with low capture cross-sections, including, for example, a carbon neutron reflector surrounding the sleeve. With the design presented here, the background neutrino production rate above the 3 MeV analysis cut is negligible compared to the signal anti-neutrino rate.
[…]
In the case of the ES single-flash signal, beam-off running allows measurement of environmental backgrounds above 3 MeV. This data can be collected during the 20% time period when the accelerator is down for maintenance. In the future, we may also investigate introducing beam chopping, which allows for nearly simultaneous beam-on and beam-off running.
2.2.1 The Contained Monoenergetic Photon Flux
As the neutrons moderate in the sleeve, nuclei are excited through collisions. The excited nuclei decay, producing monoenergetic photons. These are absorbed within the target and sleeve material; however, prior to absorption, the photons can potentially mix with BSM particles that can subsequently enter the detector and decay or interact inside.
The total energy of the resulting BSM particle is also monoenergetic. This energy signature allows for discovery through searches for sharp peaks in the energy measured within the detector if the BSM particle decays to electrons. Broad peaks are produced if the BSM particle interacts to produce scattered electrons. These signatures have a single electromagnetic flash and, therefore, must be differentiated from those environmental backgrounds with energies higher than the 3 MeV analysis cut that are seen in Fig. 3. However, beam-off running allows measurement of the environmental backgrounds and, if peaks are relatively narrow, the background estimate can be obtained by extrapolation from above and below the peak energies.
The set of monoenergetic peaks produced by photons in the source region is seen in Fig. 2. The highest flux is from isotopes produced by neutrons that are captured in the iron shielding that surrounds the source. We discuss design considerations for the shielding under the neutron flux, below. The second-highest peak, resulting from the breakup of deuterons in the cooling water, at 2.2 MeV, is below the 3 MeV analysis cut. Thus, at present, we do not consider possible BSM physics searches from this peak, but may do so in the future. Below these, a range of peaks extending above 10 MeV can provide progenitor photons for BSM particles. In summary, although the design has not been optimized for this physics (it is optimized for the electron antineutrino flux), this flux is sufficiently high to open doors to new physics.
2.2.2 The Contained Neutron Flux
The neutrons may also be progenitors of BSM particles that can penetrate the shielding around the source region. Neutron–dark sector couplings are motivated by the neutron beam/bottle experiment lifetime discrepancy [13, 14, 15, 16] . This allows them to disappear from the source area and reemerge in the LSC’s central and far side through $n\rightarrow n^{\prime}\rightarrow n$ transitions ($n\ldots$ neutron, $n^{\prime}\ldots$ dark sector neutron). IsoDAR has much higher sensitivity than surface experiments to this signature due to the large LSC detector and reduced backgrounds [17].
The signature for the neutron reappearance is a single flash from neutron capture in the center of the LSC detector. The 2.2 MeV capture on H is below the 3 MeV cut for analysis, but the 4.95 MeV capture on C, which occurs 1% of the time, will be observable within the analysis cut.
The most obvious background that must be controlled for this analysis is neutron punch-through from the source. IsoDAR has performed extensive GEANT-4 studies of shielding designs for Yemilab. These are presented in Ref. [18], but not discussed in detail in this preliminary design report, which discusses the elements of the design that are not Yemilab specific. (Yemilab specific designs will be collated into a separate design report.) The present design reduces neutron levels above 3 MeV reaching the outside of the Yemilab detector to just a few neutrons per year. For neutrons of any energy to penetrate from the outer wall of the detector to the fiducial volume, they must pass through 1.5 m of oil in the veto region, and 1.0 m of oil in the buffer region without capturing. If further protection from penetrating neutrons is desired, cuts can be placed in the fiducial region. In addition, we note that beam-off running allows studies of the environmental backgrounds.
2.3 Connecting the Fluxes to the IsoDAR Physics Program
The purpose of the design is to allow IsoDAR to offer a rich BSM physics program if paired with a kton-scale liquid scintillator detector like the LSC at Yemilab [2]. Here we briefly discuss some of the unique capabilities of the IsoDAR BSM physics program, dividing the discussion into three categories: new particles, new symmetries, and new interactions. Details of these example analyses are beyond the scope of this design report, but can be found in the references.
2.3.1 New Particles
Figure 4: An example of a potential 3+2 signal in IsoDAR, orange–true signature, black–IsoDAR reconstruction with uncertainty shown. The true oscillations are dampened due to limited statistics and energy smearing. The oscillation parameters chosen for this example are allowed by today’s experiments [19].
Example 1: Additional neutrinos—
Many new physics models introduce additional neutrinos beyond the three Standard Model (SM) flavors that have feeble or even zero SM couplings, though they may interact through new BSM couplings [20]. For light new neutrinos (with masses less than 10 MeV) it is feasible to design sensitive neutrino oscillation experiments. The probability that a neutrino interacts through the SM will vary with $L/E$, where $L$ is the distance traveled, and $E$ is the neutrino energy. The signature in the case of IsoDAR is a disappearance pattern of the $\bar{\nu}_{e}$ flux as a function of $L/E$ that will be distinct depending on the model. For example, Fig. 4 shows the complex flux disappearance pattern that arises for the case where two additional light, non-interacting neutrinos are added to the three SM neutrinos, hence called “3+2,” for an example set of oscillation parameters allowed by today’s experiments [19]. Table 1 gives the assumptions for the IsoDAR measurements in Fig. 4. The fine measurement capability in $L/E$ will allow the observed data to precisely constrain the oscillation pattern [i.e., separating (3+1) versus (3+2)], greatly narrowing the theoretical models that can produce an effect.
[…]